OpenStudy (anonymous):

Define a variable , write an equation,and solve each problem. Then check your solution. 1. Six times a number is -96. find the number 2.One fourth of a number is -16.325 . What is the number

OpenStudy (anonymous):

1.6x=-96

OpenStudy (anonymous):

6x=-96 => x=-96/6 => x=-16 (1/4)*x=-131/8 => x=(-131/8)*4 => x=-131/2

OpenStudy (anonymous):

2.(1/4)x=-16.325